a circle has an equation of 2(x-2)^2+2y^2=2. Find the center (h,k) and radius r and graph the circle. Find the intercepts if any,of the graph.
Answer by lwsshak3(11628)

a circle has an equation of 2(x-2)^2+2y^2=2. Find the center (h,k) and radius r and graph the circle. Find the intercepts if any,of the graph.
***
2(x-2)^2+2y^2=2
(x-2)^2+y^2=1
center:(2,0)
radius=1
x-intercepts
set y=0
(x-2)^2=1
x-2=±√1=±1
x=2±1
x=3, 1
y-intercepts: none

see graph below:
y=(1-(x-2)^2)^.5